OpenStudy (anonymous):

Could someone please explain how to do this step-by-step? Let L be the line with parametric equations x = −6−3t y = −5−2t z = −2−t Find the vector equation for a line that passes through the point P=P(9, −3, −4) and intersects L at a point that is distance 4 from the point Q=Q(−6, −5, −2).

OpenStudy (e.mccormick):

Not quite to this in my Lin Alg class. But I know the concept. If you are doing it through Lin Alg, this seems to discuss it pretty well: http://math.stackexchange.com/questions/348819/find-the-vector-parametric-equation-of-the-line-through-a-and-b

OpenStudy (e.mccormick):

Sorry I can't help more. Still learning that one myself!

OpenStudy (anonymous):

I'm pretty OK with the stuff in that link, but I'm having trouble with the "point distance 4 from the point Q=Q(-6,-5,-2)" - I can't figure out how to find that point (or if that's what I'm supposed to be doing first). I assumed you would add 4 times the unit vector to Q and go from there, but that doesn't seem to be right :( Thanks for the reply!

OpenStudy (e.mccormick):

Yah, that was what stumped me right off. Seeing it to a specific point, so it has to be a specific line, not just any particular line.

OpenStudy (anonymous):

I would figure out when the distance between \(L\) and \(Q\) is 4.

OpenStudy (anonymous):

\[ 4^2=(L-Q) \cdot (L-Q) \]

OpenStudy (anonymous):

That will give you at most two \(t\) values. In this case it is supposed to only give you one though. \(L(t)\) will give you the point that is passes through. With two points it is easy to find the vector equation.

OpenStudy (anonymous):

Awesome, thank you!